The number of tickets sold each day for an upcoming performance of Handel's Messiah given by N(x)=-0.5x^2+15x+12, where x is the number of days since the concert was first announced. When will daily ticket sales peak and how many tickets will be sold that day?
Answer by mananth(16949)

N(x)=-0.5x^2+15x+12
this is an inverted parabola since coefficient of x^2 is negative
x coordinate of vertex = -b/2a
b=15, a=-0.5
-b/2a= -15/2(-0.5)= -15/-1 = 15
so sales will peak on the 15th day.
plug value of x in the equation to get the y coordinate of the vertex
((-0.5) * (15^2)) + (15 * 15) + 12 = 124.5
maximum sales will be 124 on the 15th day